# PTR: A Pre-trained Language Model for Trajectory Recovery

### v1

@@ Abstract.
Spatiotemporal trajectory data is vital for web-of-things services and is extensively collected and analyzed by web-based hardware and platforms. However, issues such as service interruptions and network instability often lead to sparsely recorded trajectories, resulting in a loss of detailed movement data. As a result, recovering these trajectories to restore missing information becomes essential. Despite progress, several challenges remain unresolved. First, the lack of large-scale dense trajectory data hampers the performance of existing deep learning methods, which rely heavily on abundant data for supervised training. Second, current methods struggle to generalize across sparse trajectories with varying sampling intervals, necessitating separate re-training for each interval and increasing computational costs. Third, external factors crucial for the recovery of missing points are not fully incorporated.
To address these challenges, we propose a framework called PTR. This framework mitigates the issue of limited dense trajectory data by leveraging the capabilities of pre-trained language models (PLMs). PTR incorporates an explicit trajectory prompt and is trained on datasets with multiple sampling intervals, enabling it to generalize effectively across different intervals in sparse trajectories. To capture external factors, we introduce an implicit trajectory prompt that models road conditions, providing richer information for recovering missing points. Additionally, we present a trajectory embedder that encodes trajectory points and transforms the embeddings of both observed and missing points into a format comprehensible to PLMs. Experimental results on two public trajectory datasets with three sampling intervals demonstrate the efficacy and scalability of PTR.
@@ 1. Introduction
A trajectory is represented as a sequence of timestamped locations that depict the movement of individuals and vehicles. With the advances in web-of-things services like ride-hailing and sharing bikes and web-based platforms like Google Maps and Mapbox in recent years, massive trajectory data is collected and analyzed. Such data plays a pivotal role in various applications, including urban planning (Wang et al., 2024), traffic management (Rehena and Janssen, 2018; Wang et al., 2021), and personalized location services (Zhao et al., 2020; Wang et al., 2020).
Collecting trajectory data involves the coordination of multiple devices (Chen et al., 2024). The process begins with continuously recorded GPS coordinates from devices like smartphones or IoT sensors at regular intervals. These data points are then sent to central servers or databases for storage and analysis. However, factors such as network transmission instability, device power-saving settings, and poor coverage in high-traffic or remote areas will result in the loss of trajectory points, making the collected data sparse. Such sparse trajectories failed to accurately describe movement behavior and route choices, thus limiting their effectiveness in trajectory-based applications. To mitigate this problem, it is essential to recover the missing points from sparse trajectories which improves the integrity of trajectory information. Such task is usually termed trajectory recovery.
Many works have been proposed to tackle trajectory recovery task, which can be divided into two categories: free space trajectory recovery (Chen et al., 2011, 2023a) and map-matched trajectory recovery (Wei et al., 2024; Chen et al., 2023b). Free space trajectory recovery aims to directly restore the coordinates of missing trajectory points but cannot ensure that the recovered trajectory is constrained to the road network. Typically, an additional map-matching process is required before these trajectories can be used in navigation-based applications. Map-matched trajectory recovery focuses on directly recovering trajectories on the road network. In this paper, we focus on the latter setting because it can be directly utilized by downstream tasks without the need for further processing. Despite the progress made by existing methods on map-matched trajectory recovery, they still face some challenges as listed below.
1) The availability of dense trajectories is often limited. Deep learning-based trajectory recovery models are inherently data-driven, requiring a substantial number of dense trajectories as labels for supervised training. However, due to the aforementioned factors, most collected and stored trajectories are sparse, resulting in a limited availability of large-scale dense trajectory data. As a result, the lack of sufficient dense trajectories for training makes existing deep learning models prone to overfitting and leads to diminished performance.
2) Existing models struggle to generalize to sparse trajectories with varying sampling intervals. Sparse trajectories often exhibit inconsistent sampling intervals due to network instability. However, most trajectory recovery models are trained on sparse trajectories with fixed sampling intervals, limiting their ability to generalize to trajectories with varying intervals. As a result, when confronted with trajectories that have different sampling intervals, these models require separate training for each interval to recover trajectories effectively, leading to significantly increased training costs.
Figure 1. An example of road network evolution process.
3) External factors related to the recovery of missing points are not incorporated. External factors, such as road conditions, provide valuable insights that can enhance trajectory recovery. For instance, in the scenario illustrated in Figure 1, given the observed points $p_{1}$ and $p_{4}$ and the road conditions at time $t_{1}$, it is more likely that the user took route $R_{1}$ rather than $R_{2}$. Additionally, based on the road conditions at times $t_{2}$ and $t_{3}$, as shown in Figures 1 (b) and (c), it becomes easier to infer that the user first accelerated and then decelerated between $p_{1}$ and $p_{4}$. However, existing methods only consider information from the known trajectory points, overlooking the significance of external factors.
To address the aforementioned challenges, we propose a Pre-trained Language Model for Trajectory Recovery (PTR). By leveraging a pre-trained language model (PLM) that has been trained on large-scale corpus datasets and fine-tuning it with a limited set of dense trajectories, PTR can effectively recover trajectories even when only scarce dense trajectory data is available. PTR incorporates two key components to generalize to sparse trajectories with varying sampling intervals: first, an explicit trajectory prompt, which uses natural language to describe both the trajectory recovery task and its characteristics. This activates the PLM’s ability to recover trajectories and recognize the trajectory’s sampling interval. Second, a joint training strategy, where PTR is trained on trajectory data with multiple sampling intervals, further enhancing its generalization capabilities. To account for external factors, we introduce an implicit trajectory prompt, which integrates rich contextual information, such as road conditions, to provide insights on missing points. Lastly, we propose a trajectory embedder that encodes both observed and missing trajectory points, transforming them into representations interpretable by PLMs. This allows the PLM to capture the spatiotemporal correlations within the trajectory.
@@
  • We introduce PTR, a novel framework for trajectory recovery that utilizes PLMs to efficiently recover trajectories using a limited set of dense trajectories.
  • We propose explicit trajectory prompts that enable the PLM to accurately interpret the trajectory recovery task and determine the sampling intervals of the trajectories. Combined with a joint training strategy on trajectories with various sampling intervals, PTR generalizes effectively to sparse trajectories across different intervals.
  • We present an implicit trajectory prompt to model external factors influencing missing trajectory points and a trajectory embedder to transform the trajectory representation into a form that the PLM can process.
  • We conduct extensive experiments on two real-world trajectory datasets, each with three different sampling intervals. The results show that our proposed model achieves state-of-the-art performance in both accuracy and efficiency.
@@ 2. Related Work
2.1. Free-space Trajectory Recovery
Trajectory recovery in free space aims to directly restore the GPS coordinates of missing points. Early approaches relied on predefined rules to model mobile objects (Chen et al., 2016; Su et al., 2013). For example, Linear (Hoteit et al., 2014) assumes that vehicle movement follows uniform linear motion and employs linear interpolation for trajectory recovery. MPR (Chen et al., 2011) assumes vehicles travel along the most frequently used routes, utilizing historical trajectory data to determine route frequencies for recovery. Other studies (Banerjee et al., 2014; Zheng et al., 2012) use Markov models to capture spatial transitions between trajectory points. However, these methods are limited to modeling low-order transitions and struggle to capture the global spatial-temporal dependencies of trajectories.
Recent deep learning-based approaches have provided more effective solutions for trajectory recovery (Xu et al., 2017, 2022). DHTR (Wang et al., 2019) employs a GRU-based seq2seq model to analyze user route transitions, integrating Kalman filtering to improve recovery accuracy. AttnMove (Xia et al., 2021) uses attention mechanisms and LSTM to capture overall human movement preferences. PeriodicMove (Sun et al., 2021) models each user trajectory as a graph, employing graph neural networks to efficiently capture user preferences by leveraging historical route transitions. TERI (Chen et al., 2023a) introduces a two-stage framework based on transformers to recover trajectories with unequal spacing, while TrajBERT (Si et al., 2023) utilizes the BERT (Devlin et al., 2018) framework to recover implicit trajectories, focusing on coarser-grained outputs such as area or Wi-Fi IDs. Although these methods show promising results, they require an additional map-matching process to align the recovered trajectories with the road network before they can be applied in navigation systems. This two-stage process leads to error accumulation and introduces significant time overhead, as the map-matching step is time-consuming.
2.2. Map-matched Trajectory Recovery
Compared with the above methods, map-matched trajectory recovery incorporates the road network as input and aims to recover the trajectory directly onto the road. To this end, MTrajRec (Ren et al., 2021) first represents the trajectory point with a road segment and moving rate, utilizing the seq2seq-based multi-task framework to capture the spatiotemporal correlations within the trajectory. Following this idea, RNTrajRec (Chen et al., 2023b) further models the relationship between trajectories and road networks, designing a spatial-temporal transformer architecture to encode sparse trajectories. MM-STGED (Wei et al., 2024) models sparse trajectories from a graph perspective, considering both micro and macro semantic information of trajectories.
Although these methods demonstrate promising progress in trajectory recovery, their performance are still hindered by the poor availability of large-scale dense trajectories and limited in generalization to different sampling intervals discussed in Section 1.
@@ 3.1. Definitions
By employing the map-matching algorithm, a trajectory $\mathcal{T}$ can be projected onto the road network and return the map-matched trajectory $\mathcal{T}_{m}$. As a result, each point of the map-matched trajectory aligns accurately with a particular road. The map-matched trajectory is represented as $\mathcal{T}_{m}=\langle q_{1},\cdots,q_{|\mathcal{T}_{m}|}\rangle$, where each $q_{j}=(e_{j},r_{j},t_{j})$ signifies the vehicle’s location at time $t_{j},j\in\{1,\cdots,|\mathcal{T}_{m}|\}$. Here, $e\in\mathcal{V}$ is the matched road segment, and $r$ is the moving ratio, which quantifies the proportion of the distance traveled from the starting point of a road segment in relation to the total length of that road segment.
@@
Figure 3. The framework of PTR, consists of Trajectory Prompt, Trajectory Embedder, and PLM-based Trajectory Encoder.
@@ 4.1. Overall Pipeline
In this paper, we present PTR, which fine-tunes a PLM pre-trained on a large-scale corpus to efficiently recover trajectories with limited dense data. The framework of PTR, shown in Figure 3, comprises three main components: trajectory prompts, a trajectory embedder, and a PLM-based trajectory encoder.
Trajectory prompts offer essential information from two key components. First, explicit trajectory prompts, expressed in natural language, outline the trajectory recovery task and key trajectory characteristics. These prompts activate the PLM’s ability to recover trajectories and identify varying sampling intervals in sparse trajectories. With joint training on trajectories across different intervals, PTR achieves robust generalization. Second, implicit trajectory prompts model road conditions, providing crucial external factors of missing points.
The trajectory embedder effectively enhances the PLM’s understanding of trajectories. It begins by normalizing sparse trajectories with varying sampling intervals $\mu$ into a uniform interval $\epsilon$. Next, each trajectory point, including both observed and missing points, is encoded into an embedding that captures both local and global semantics. Finally, the embedding is transformed into a format understood with the PLM.
In the PLM-based trajectory encoder, most of the encoder’s parameters are frozen to preserve the pre-trained knowledge, while the multi-head attention layer remains trainable, allowing the model to capture the spatiotemporal correlations within the trajectory. Finally, the output layer predicts the road segment $e$ and the moving ratio $r$ for the recovered map-matched trajectory at each step.
4.2. Trajectory Prompts
To enable PLMs to recover trajectories with different intervals and efficiently model external factors for missing trajectory points, we introduce explicit and implicit trajectory prompts.
4.2.1. Explicit Trajectory Prompt
The explicit trajectory prompt offers a structured textual description at both the task and trajectory levels, enabling PTR to adapt to varying sampling intervals. At the task level, it includes the following components:
@@ 4.2.1. Explicit Trajectory Prompt
At the trajectory level, the explicit trajectory prompt provides detailed characteristics of the trajectory, consisting of:
@@
The ¡Time Part¿ provides the trajectory’s specific start and end times, helping the PLM understand the duration and potential time patterns, such as morning or evening peaks. The ¡Movement Part¿ supports the PLM in inferring the trajectory’s movement.
After obtaining the prompt for each part, we use the tokenizer and word token embedding in PLMs to convert text into embeddings. The embedding of ¡Task Part¿, ¡Target Part¿, ¡Content Part¿, ¡Time Part¿, and ¡Movement Part¿ are denoted by $\textbf{H}_{ta}$, $\textbf{H}_{tar}$, $\textbf{H}_{c}$, $\textbf{H}_{t}$, and $\textbf{H}_{m}$. These embeddings are then concatenated to form the overall explicit trajectory prompt embedding.
@@
4.2.2. Implicit Trajectory Prompt
We incorporate road conditions as external factors for missing trajectory points, as they provide insights into both the surrounding environment and the object’s movement, which are all helpful for trajectory recovery. For instance, vehicles typically slow down in congested areas and accelerate in less congested regions. Given the complexity and variability of real-world road conditions, which are difficult to express in natural language, we represent them as an implicit trajectory prompt.
To obtain the road conditions for each trajectory point, we first calculate the average road conditions across all areas and time intervals, then extract the relevant information for each point. Specifically, we divide the area of interest into a spatial grid of $I\times J$ cells and partition the time into $T$ slices. For each grid cell, we compute the total traffic, resulting in a regional flow matrix $\textbf{RC}\in\mathbb{R}^{I\times J\times T}$. To capture the spatiotemporal correlations among different regions, we apply a 2D convolution (2D CNN) along the spatial dimensions and a 1D CNN along the temporal dimension, yielding the road condition representation $\textbf{H}_{rc}\in\mathbb{R}^{I\times J\times T\times F}$, where $F$ is the number of features. Formally:

### v2

# PLMTrajRec: A Scalable and Generalizable Trajectory Recovery Method with Pre-trained Language Models

@@ Abstract.
Spatiotemporal trajectory data is crucial for various applications. However, issues such as device malfunctions and network instability often cause sparse trajectories, leading to lost detailed movement information. Recovering the missing points in sparse trajectories to restore the detailed information is thus essential. Despite recent progress, several challenges remain. First, the lack of large-scale dense trajectory data makes it difficult to train a trajectory recovery model from scratch. Second, the varying spatiotemporal correlations in sparse trajectories make it hard to generalize recovery across different sampling intervals. Third, the lack of location information complicates the extraction of road conditions for missing points.
To address these challenges, we propose a novel trajectory recovery model called PLMTrajRec. It leverages the scalability of a pre-trained language model (PLM) and can be fine-tuned with only a limited set of dense trajectories. To handle different sampling intervals in sparse trajectories, we first convert each trajectory’s sampling interval and movement features into natural language representations, allowing the PLM to recognize its interval. We then introduce a trajectory encoder to unify trajectories of varying intervals into a single interval and capture their spatiotemporal relationships. To obtain road conditions for missing points, we propose an area flow-guided implicit trajectory prompt, which models road conditions by collecting traffic flows in each region. We also introduce a road condition passing mechanism that uses observed points’ road conditions to infer those of the missing points. Experiments on two public trajectory datasets with three sampling intervals each demonstrate the effectiveness, scalability, and generalization ability of PLMTrajRec.
Trajectory recovery, Pre-trained language model, Road network
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: KDD 2025; August 03–07, 2025; Toronto, Canadaisbn: 978-1-4503-XXXX-X/XX/XX
@@ 1. Introduction
A trajectory is a sequence of timestamped locations that describe the movement of individuals or vehicles, represented as $\mathcal{T}=\langle p_{1},\dots,p_{|\mathcal{T}|}\rangle$ of length $|\mathcal{T}|$, where $p_{i}=(lat_{i},lng_{i},t_{i})$ denotes the latitude, longitude, and timestamp of the $i$-th point. The sampling interval of $\mathcal{T}$ is $\mu=t_{i}-t_{i-1},\forall i\in\{2,\dots,|\mathcal{T}|\}$. With the advances in GPS devices and geopositioning technologies, a massive amount of trajectory data has been collected. These data play a pivotal role in various applications, such as urban planning (Wang et al., 2024), traffic management (Rehena and Janssen, 2018; Wang et al., 2021), and personalized location services (Zhao et al., 2020; Wang et al., 2020). However, factors such as network instability, device malfunctions, or cost-saving settings often lead to missing trajectory points, making the data sparse with large sampling intervals (Ren et al., 2021; Zhao et al., 2019). Sparse trajectories fail to accurately reflect movement behavior and route choices, limiting their usefulness. To address this issue, recovering the missing points in sparse trajectories is crucial for preserving trajectory completeness—a task usually referred to as trajectory recovery.
Many works have been proposed to tackle trajectory recovery, broadly categorized into free space trajectory recovery (Chen et al., 2011, 2023a) and map-matched trajectory recovery (Wei et al., 2024; Chen et al., 2023b). Free space methods directly predict the coordinates of missing points but do not ensure alignment with the road network, typically requiring a subsequent map-matching step. Map-matched methods, by contrast, aim to recover missing points directly on the road network, making them more convenient for downstream applications. Despite the progress of existing methods, map-matched trajectory recovery still faces the following challenges:
1) Limited dense trajectory data. Trajectory recovery models are data-driven and typically require large-scale pairs of sparse and dense trajectories. However, because of cost-saving settings and device malfunctions, most collected trajectories remain sparse. The limited availability of dense trajectory data makes existing models prone to overfitting and hampers their performance.
2) Difficulty in generalizing to varying sampling intervals. Due to network instability and device malfunctions, a sparse trajectory dataset often contains trajectories with a mixture of different sampling intervals (Wang et al., 2021). Existing works (Wei et al., 2024; Chen et al., 2023b; Ren et al., 2021) usually treat all intervals in the same way, overlooking the varying spatiotemporal correlations of sparse trajectories entangled with multiple sampling intervals (Chen et al., 2024; Che et al., 2018). As a result, when encountering trajectories with unseen sampling intervals, these models often require retraining, which increases costs.
Figure 1. Impact of road conditions on route selection and movement patterns.
3) Non-trivial extraction of dynamic road conditions for missing points. Road conditions are key to understanding movement patterns. Existing work (Wei et al., 2024) incorporates road conditions around observed points but ignores those at missing points. For example, in Figure 1(b) and (c), knowing the conditions at $p_{2}$ and $p_{3}$ reveals that the user is gradually decelerating on $R_{1}$ due to congestion. If we only consider conditions at $p_{1}$ and $p_{4}$, we might infer that the user takes $R_{1}$ instead of $R_{2}$ (Figure 1(a)), but not the finer details of the movement. Since the exact locations of missing points are unknown, their road conditions cannot be directly extracted.
To tackle these challenges, we propose Pre-trained Language Model for Trajectory Recovery (PLMTrajRec). We leverage a pre-trained language model (PLM) that has already learned useful representations from large-scale corpus data and fine-tune it using only a small number of dense trajectories. This enables PLMTrajRec to recover missing points even when dense trajectories are scarce (Challenge 1). To handle sparse trajectories with varying sampling intervals (Challenge 2), we introduce an interval and feature-guided (IF-guided) explicit trajectory prompt. It encodes both sampling intervals and movement features into the PLM through natural language, helping the model identify these intervals and capture key trajectory characteristics. We then propose an interval-aware trajectory embedder to standardize different sampling intervals and learn their spatiotemporal correlations. To infer road conditions for missing points (Challenge 3), we design an area flow-guided (AF-guided) implicit trajectory prompt that gathers traffic flows in each region. We also present a road condition passing mechanism that uses road conditions from nearby observed points to estimate those of the missing points.
@@
  • We introduce PLMTrajRec, a trajectory recovery model with strong scalability. By leveraging a pre-trained language model, PLMTrajRec can be fine-tuned on a small set of dense trajectories while still achieving effective recovery.
  • We propose an IF-guided explicit trajectory prompt and an interval-aware trajectory embedder to help PLMTrajRec generalize to sparse trajectories with various sampling intervals. The prompt encodes intervals and movement features into the PLM, and the embedder unifies intervals and learns spatiotemporal correlations.
  • We design an AF-guided implicit trajectory prompt to capture road conditions by aggregating traffic flows, and a road condition passing mechanism to infer missing-point conditions from nearby observations.
  • We conduct extensive experiments on two real-world datasets, each with three sampling intervals, showing that PLMTrajRec achieves state-of-the-art performance in effectiveness, scalability, and generalizability.
@@ 2. Related Work
Free-space Trajectory Recovery. Trajectory recovery in free space aims to directly restore missing GPS coordinates. Early methods relied on predefined rules to model mobile objects (Chen et al., 2016; Su et al., 2013), such as the vehicle moving with uniform linear motion (Hoteit et al., 2014) or taking the most popular route (Chen et al., 2011). Other approaches (Banerjee et al., 2014; Zheng et al., 2012) use Markov models to capture spatial transitions. However, these methods are limited to modeling low-order transitions and fail to capture the global spatial-temporal dependencies essential for accurate trajectory recovery. Recent deep learning-based methods have provided more effective solutions (Xu et al., 2017, 2022). DHTR (Wang et al., 2019) employs a GRU-based seq2seq model to analyze user route transitions, integrating Kalman filtering to improve accuracy. AttnMove (Xia et al., 2021) leverages attention mechanisms and LSTM to capture movement preferences, PeriodicMove (Sun et al., 2021) models trajectories as graphs using GNNs to extract user patterns. TERI (Chen et al., 2023a) employs a transformer-based two-stage framework for irregular interval trajectory recovery, and TrajBERT (Si et al., 2023) adapts BERT (Devlin et al., 2018) for implicit trajectory recovery at a coarser granularity. Despite their effectiveness, these methods require an additional map-matching step to align recovered trajectories with road networks before they can be applied in navigation systems. This two-stage process leads to error accumulation and significant time overhead due to the computational cost of map-matching.
Map-matched Trajectory Recovery. Compared with the above methods, map-matched trajectory recovery incorporates the road network as input and aims to recover the trajectory directly onto the road. To this end, MTrajRec (Ren et al., 2021) first represents the trajectory point with a road segment and moving rate, utilizing the seq2seq-based multi-task framework to capture the spatiotemporal correlations within the trajectory. Following this idea, RNTrajRec (Chen et al., 2023b) models the relationship between trajectories and road networks, designing a spatial-temporal transformer architecture to encode sparse trajectories. LightTR (Liu et al., 2024) proposes a lightweight trajectory recovery framework based on federated learning. MM-STGED (Wei et al., 2024) models sparse trajectories from a graph perspective, considering both micro and macro semantic information of trajectories. Although they demonstrate promising progress in trajectory recovery, their performance is still hindered by the poor availability of large-scale dense trajectories and limited in generalization to different sampling intervals discussed in Section 1.
@@ 3.1. Definitions
By employing the map-matching algorithm, a trajectory $\mathcal{T}$ can be projected onto the road network and return the map-matched trajectory $\mathcal{T}_{m}$. As a result, each point of the map-matched trajectory aligns accurately with a particular road. The map-matched trajectory is represented as $\mathcal{T}_{m}=\langle q_{1},\cdots,q_{|\mathcal{T}_{m}|}\rangle$, where each $q_{j}=(e_{j},r_{j},t_{j})$ signifies the vehicle’s location at time $t_{j},j\in\{1,\cdots,|\mathcal{T}_{m}|\}$. Here, $e\in\mathcal{E}$ is the matched road segment, and $r$ is the moving ratio, which quantifies the proportion of the distance traveled from the starting point of a road segment in relation to the total length of that road segment.
@@
Figure 3. The framework of PLMTrajRec, consists of Dual Trajectory Prompts, Interval-aware Trajectory Embedder, and PLM-based Trajectory Encoder.
@@ 4.1. Overall Pipeline
In this paper, we present both scalable and generalizable trajectory recovery model, PLMTrajRec, by fine-tuning a PLM that is pre-trained on a large-scale corpus with limited dense trajectory data. The framework of PLMTrajRec, shown in Figure 3, comprises three main components: dual trajectory prompts, an interval-aware trajectory embedder, and a PLM-based trajectory encoder.
The dual trajectory prompts provide essential information through two key components. First, the interval and feature-guided (IF-guided) explicit trajectory prompt, expressed in natural language, injects the sampling interval of sparse trajectories and their movement features into PLM. This enhances the model’s ability to recognize the sampling intervals of each sparse trajectory and capture trajectory characteristics. Second, area flow-guided (AF-guided) implicit trajectory prompts model road conditions, offering crucial contextual information to improve the recovery of missing trajectory points.
The interval-aware trajectory embedder first normalizes sparse trajectories with varying sampling intervals $\mu$ into a unified interval $\epsilon$, effectively handling the diverse spatiotemporal correlations associated with different sampling intervals and enhancing model generalization. Subsequently, each trajectory point, including both observed and missing points, is encoded into an embedding that the PLM can process.
In the PLM-based trajectory encoder, pre-trained BERT is employed to capture information bi-directionly, and most of the parameters are frozen to preserve the pre-trained knowledge, while the multi-head attention layer remains trainable, allowing the model to capture the spatiotemporal correlations within the trajectory. Finally, the output layer predicts the road segment $e$ and the moving ratio $r$ for the recovered map-matched trajectory at each step.
4.2. Dual Trajectory Prompts
To enable PLM to recover trajectories with different intervals and effectively model road conditions for missing trajectory points, we introduce dual trajectory prompts, including an interval and feature-guided (IF-guided) explicit trajectory prompt and an area flow-guided (AF-guided) implicit trajectory prompt.
4.2.1. IF-guided Explicit Trajectory Prompt
The IF-guided explicit trajectory prompt offers a structured textual description of the sampling intervals and movement features of sparse trajectories, enabling PLMTrajRec to identify varying sampling intervals and capture trajectory features. The prompts related to the sampling interval include the following components:
@@ 4.2.1. IF-guided Explicit Trajectory Prompt
The prompts related to the movement features include the following contents:
@@
The ¡Time Part¿ provides the trajectory’s specific start and end times, helping the PLM understand the duration and potential time patterns, such as morning or evening peaks. The ¡Movement Part¿ supports the PLM in inferring the trajectory’s movement. We give a detailed example of the IF-guided explicit trajectory prompt in Appendix C.
After obtaining the prompt for each part, we use the tokenizer and word token embedding in PLMs to convert text into embeddings. The embedding of ¡Task Part¿, ¡Target Part¿, ¡Content Part¿, ¡Time Part¿, and ¡Movement Part¿ are denoted by $\textbf{H}_{ta}$, $\textbf{H}_{tar}$, $\textbf{H}_{c}$, $\textbf{H}_{t}$, and $\textbf{H}_{m}$. These embeddings are then concatenated to form the overall IF-guided explicit trajectory prompt embedding.
@@
4.2.2. AF-guided Implicit Trajectory Prompt
Road conditions can provide insights into both the surrounding environment and the object’s movement, which are helpful for trajectory recovery. For instance, vehicles typically slow down in congested areas and accelerate in less congested regions. Given the complexity and variability of real-world road conditions are difficult to express in natural language, we represent them as the implicit trajectory prompt.
To obtain the road conditions for each trajectory point, we first calculate the average road conditions across all areas and time intervals, then extract the relevant information for each point. Specifically, we divide the area of interest into a spatial grid of $I\times J$ cells and partition the time into $T$ slices. For each grid cell, we compute the total traffic, resulting in a regional flow matrix $\textbf{RC}\in\mathbb{R}^{I\times J\times T}$. Each element in RC represents the number of traffic flows at region $i,j$ at time $t$. To capture the spatiotemporal correlations among different regions, we apply a 2D convolution (2D CNN) along the spatial dimension and a 1D CNN along the temporal dimension, yielding the road condition representation $\textbf{H}_{rc}\in\mathbb{R}^{I\times J\times T\times F}$, where $F$ is the number of features. Formally: